OpenStudy (anonymous):

What is the height of this triangle? A. 4.5 ft B. 9 ft C. 49 ft D. 882 ft

OpenStudy (anonymous):

what triangle?

OpenStudy (cicilybailey14):

I see no triangle.

OpenStudy (anonymous):

Triangle?

OpenStudy (anonymous):

yey me either

OpenStudy (anonymous):



OpenStudy (anonymous):

o there it is .............

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

u gotz to learn to be patient

sammixboo (sammixboo):

Ok, so what is the formula to find the area of a triangle?

OpenStudy (anonymous):

Well it WOULD deffently NOT be C or D .......... what do u think it is ?

OpenStudy (anonymous):

BxH

OpenStudy (anonymous):

Master mind it is 9ft

OpenStudy (anonymous):

But.

OpenStudy (anonymous):

you need to understand why

sammixboo (sammixboo):

Hold on a sec, haha

OpenStudy (anonymous):

If the formula is bxh divided by two

OpenStudy (anonymous):

And if the total area is 63

OpenStudy (anonymous):

eh i asked elleven questions today and i got 3 wrong

OpenStudy (anonymous):

you multiply 63 by 2

sammixboo (sammixboo):

The formula to find your height is \(\rm \color{20bd23}{h~=~\dfrac{2\times A}{b}}\) Plug in the info given in the pictur \(\rm \color{20bd23}{h~=~\dfrac{2\times 63}{14}}\) What is 2 x 63?

OpenStudy (anonymous):

that took alot more time to type than it should have lol

OpenStudy (anonymous):

and i already answerd it it was 9

sammixboo (sammixboo):

I am having internet problems :(

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

yep